# Escaping Plato’s Cave: JAM for Aligning Independently Trained Vision and Language Models

### v1

@@ Abstract
Independently trained vision and language models inhabit disjoint representational spaces, shaped by their respective modalities, objectives, and architectures. Yet an emerging hypothesis — the Platonic Representation Hypothesis — suggests that such models may nonetheless converge toward a shared statistical model of reality. This compatibility, if it exists, raises a fundamental question: can we move beyond post-hoc statistical detection of alignment and explicitly optimize for it between such disjoint representations? We cast this Platonic alignment problem as a multi-objective optimization task: preserve each modality’s native structure while aligning for mutual coherence. We introduce the Joint Autoencoder Modulator (JAM) framework that jointly trains modality-specific autoencoders on the latent representations of pre-trained single modality models, encouraging alignment through both reconstruction and cross-modal objectives. By analogy, this framework serves as a method to escape Plato’s Cave, enabling the emergence of shared structure from disjoint inputs. We evaluate this framework across three critical design axes: (i) the alignment objective — comparing contrastive loss (Con), its hard-negative variant (NegCon), and our Spread loss; (ii) the layer depth at which alignment is most effective; and (iii) the impact of foundation model scale on representational convergence. Our findings show that our lightweight Pareto-efficient framework reliably induces alignment, even across frozen, independently trained representations, offering both theoretical insight and practical pathways for transforming generalist unimodal foundations into specialist multimodal models.
@@ 1 Introduction
Neural networks trained on different modalities, datasets, and objectives typically inhabit disjoint representational spaces. Yet an emerging hypothesis suggests that these models — despite having no shared supervision, architecture, or training regime — may still converge toward a common statistical model of reality. This hypothesis is known as the Platonic Representation Hypothesis (PRH) [1], whose naming is inspired by Plato’s Allegory of the Cave [2], which describes individuals who observe only shadows of objects cast on the wall of a cave and mistake these projections for the entirety of reality. In our context, training data and unimodal representations are the “shadows” or “projections” of reality. The true structure, or Platonic representation, lies behind the wall: a latent alignment of semantics that multiple models might uncover despite training in isolation.
This hypothesis has been discussed under several philosophical and empirical lenses, including convergent realism in the philosophy of science and the Anna Karenina scenario [3] in representation learning, which suggests that all well-performing models may ultimately resemble each other. However, most evidence to date remains observational, relying on statistical metrics that infer compatibility. These methods quantify correlations but do not suggest practical recipies for creating multimodal models from aligning unimodal ones. In this work, we develop a method to escape Plato’s Cave, enabling two independently trained models with different modalities to work together.
@@ 1.1 Escaping Plato’s Cave (Platonic Alignment)
Inspired by the Platonic Representation Hypothesis [1], we ask the question: “can unimodal models, trained independently on distinct modalities and objectives, nonetheless encode structurally compatible representations that can be discovered and aligned for post hoc?” We frame this goal as Platonic Alignment. To investigate this question, we develop the Joint Autoencoder Modulator (JAM) — a lightweight modality-specific modules that aligns frozen unimodally trained language and vision models via coordinated reconstruction and alignment objectives. Treating unimodal embeddings as fixed inputs, JAM jointly trains two modality-specific autoencoders that retain each modality’s native structure while encouraging representational coherence in their shared bottleneck. In doing so, the reconstruction objective anchors modality-specific information, while the alignment objective facilitates the emergence of shared conceptual space — a computational escape from Plato’s Cave. Our work makes the following contributions:
  • We introduce and compare multimodal alignment objectives – classic contrastive learning (Con), its hard-negative variant (NegCon), and our proposed Spread loss, which encourages structured alignment by leveraging the notion of shared context across texts corresponding to image. Notably, Spread loss consistently outperforms the other objectives, demonstrating the benefit of leveraging contextual structure for fine-grained alignment.
  • We utilized curriculum-based approach as a probing strategy to examine how layer depth (early, mid, late) of pretrained encoders impacts alignment behavior and supervision needs.
  • We analyze the role of model scale by comparing alignment performance across language backbones of varying parameter counts, revealing how scale interacts with limited supervision.
@@ 1.2 Related Work
Testing for Alignment. Prior work in this direction has been largely diagnostic focusing on evaluating alignment between frozen features: measures such as centered kernel alignment (CKA) [4], variants of CCA (SVCCA [5], projection-weighted CCA [6]), and nearest-neighbor metrics [7, 1]. Other approaches explore probing tasks or zero-shot transfer to assess latent compatibility across modalities [8, 9]. More recently, model-stitching frameworks have examined whether independently trained sub-networks can be functionally composed to perform new tasks [3], hinting at deeper interoperability between pretrained systems but focusing on vision domain.
@@ 2 Statistical Tests for Representation Alignment
Figure 1: Statistical Metrics for Representation Alignment: (a) Linear and Kernel (Gaussian (i.e., RBF) kernel) PCA to reduce the embeddings dimension to 50, Set CCA dimension as 50. Variation of score values are based on the utilized kernel.; (b) Variation of score values are based on the utilized kernel (Linear and RBF); (c) SVD to reduce embeddings dimension to 10, Set CCA dimension as 10.; (d) Set top k nearest neighbors = 10 (following the default setting of [1]). Across all metrics and model cases, match pairs consistently show higher alignment scores than easy non-match pairs, supporting the hypothesis that unimodal models encode shared global structure. However, hard non-match pairs exhibit similarly high scores. This indicates that while post-hoc similarity metrics may reveal coarse representational compatibility – which aligns with PRH [1] – they are insufficient for diagnosing semantic alignment at a more granular level.
Before introducing our Platonic alignment framework, we first assess the extent to which alignment between unimodal vision and language models can be detected through statistical metrics alone. We construct 3 settings differing in their underlying semantic alignment:
  • Matching pairs (true positive), where the image and text both refer to the same real-world content — they form a true positive pair sharing an instance of reality (e.g., an image of a man sitting in a cave and the caption “a man sitting in a cave”). Consistent with the Platonic Representation Hypothesis, these pairs should yield high alignment scores, reflecting shared conceptual grounding.
  • Easy non-matching pairs (easy negatives), where the image and text are drawn from unrelated and incoherent sources — such as a white noise image paired with real text, or a real image paired with syntactically valid but semantically meaningless — representing negative pairs with no shared underlying content. These pairs should yield very low alignment scores, as they lack any shared semantic or structural foundation.
  • Hard non-matching pairs (hard negatives), where the image remains fixed but the text is a hard-negative caption: a semantically similar alternative that diverges from the true caption in fine-grained details (e.g., an image of a man sitting in a cave, but the corresponding caption is "a man sitting on the moon"), thus contextually aligned but not referentially accurate. While more challenging, the PRH would still predict lower alignment scores here if the models encode referential distinctions; high scores in this case would suggest limitations of statistical metrics in capturing fine-grained semantic mismatch.
We use the 4 representation alignment metrics or scores to compare embedding spaces: Canonical Correlation Analysis (CCA) [18], Centered Kernel Alignment (CKA) [4], SVCCA [5], and CKNNA [1]. (Metrics formulation in Appendix.)
2.1 Data & Model Set-up
We use the SugarCrepe benchmark [19] — a dataset designed to test fine-grained vision-language compositionality. Given an image, the task requires selecting between the correct positive caption and a hard negative caption. These negative captions are carefully constructed to differ from the correct caption through minimal, often localized compositional changes. This setup is ideal for evaluating models that aim to resolve fine-grained semantic distinctions rather than broad categorical matches.
SugarCrepe [19] defines seven types of hard negative captions, categorized under three compositional transformation families:
  • Replace (R): Substitutes an atomic concept—attribute, object, or relation—with an incorrect alternative (e.g., “a white cave” → “a brown cave”). Includes Replace-Att (RA), Replace-Obj (RO), and Replace-Rel (RR).
  • Swap (S): Exchanges two elements of the same type, altering structure while preserving content (e.g., “a man in the cave” → “the cave in a man”). Includes Swap-Att (SA) and Swap-Obj (SO).
  • Add (A): Inserts a new, incorrect concept into the caption, introducing subtle mismatches (e.g., “a man on the moon” → “a man on the blood moon”). Includes Add-Att (AA) and Add-Obj (AO).
@@ 2.1 Data & Model Set-up
With this data, we extracted text embeddings (of all hidden states/layers) and image embeddings (CLS token embeddings of all hidden states/layers) from unimodally pretrained language and vision models, respectively. Pretrained unimodal models utilized for all the experiments are: Gemma2 (2B) [20], Llama3.2 (1B) [12], OLMo2 (7B) [21] for language models, and DINOv2 (ViT-B) [22] for vision model. With the extracted feature representations, let $D$ be the whole data, and $B$ be the batch. We construct $D$ in the nested pair format: $D=\{(v_{i},(l_{P_{i}},l_{N_{i}})\}_{i=1}^{n}$ (Refer to Table 5 for variable descriptions). $C(i)$, the set of similar context text embedding for anchor $i$, is defined as $l_{P_{i}}$ and its hard negative $l_{N_{i}}$ (i.e., $C(i)=\{l_{P_{i}},l_{N_{i}}\}$). Accordingly, $\widetilde{C}(i)=L\setminus C(i)$.
2.2 Results: Statistical tests detect alignment, but not in a fine-grained manner.
The results (Fig.1) show that all models reliably distinguish between matching pairs and easy non-matches, with statistical metrics consistently reflecting lower similarity scores for mismatched pairs. This indicates that even when trained independently, vision and language models encode representations that reflect a shared high-level structure. This observation aligns with the Platonic Representation Hypothesis[1], which suggests that different modalities—despite being trained with distinct objectives—can converge toward a common statistical model of reality, enabling coarse-grained compatibility across unimodal representations.
@@
However, we observe a compelling pattern in the hard non-match condition: statistical similarity metrics yield scores nearly indistinguishable from the scores of true matching pairs. Rather than indicating a failure of alignment, this observation reveals an important insight into the statistical tests and its interaction with the hard negatives: hard negative examples’ semantic content overlaps extensively with the positive pairs at the global contextual level, differing only in fine-grained attributes. Thus, we conjecture that the representational similarity is real and consistent with shared underlying structure, but it may not be fine-grained enough to capture discriminative semantic detail. This exposes a key limitation of statistical metrics: while they can reveal high-level compatibility, they obscure nuanced yet critical distinctions for fine-grained alignment. To move beyond this limitation, we turn to a lightweight post-hoc training approach that explicitly optimizes for semantic coherence, both globally and locally. This insight directly led to our alignment training scheme in the next section.
@@ 3 Our method: Joint Autoencoder Modulator (JAM)
3.1 Loss functions
We develop a few ideas for loss function, some of which are leveraging existing work and others developed for JAM. We introduce Contrastive and Negative Contrastive loss which are used as a baseline as well as inspire our final loss: Spread Loss functions for vision-to-language and language-to-vision. Having a set of losses designed to optimize different aspects of the alignment had a side benefit of helping us to diagnose and analyze experiments (e.g., which layer to establish alignment), as will be shown in later sections.
Contrastive Loss (Con)
@@ 3.1 Loss functions
To address this, we introduce $\mathcal{L}_{spread}$, a contrastive objective that incorporates a notion of context similarity and fine-grained differentiation. The core idea is to pull together all semantically similar captions, including the ground-truth caption and selected hard negatives, and then selectively push apart fine-grained distinctions within that similar group. This two-stage formulation allows us to preserve shared contextual alignment while enforcing discriminative power at the instance level. The idea of spread loss has been proposed in the domain of visual representation learning [29, 30], which is a variant of Supervised Contrastive (SupCon) loss [31], but tailored to tackling class collapse. We take inspiration from that work and apply this idea to the multi-modal alignment problem. [^1] We provide a parallel comparison between $\mathcal{L}_{spread}$ of multimodal representation space (i.e., ours) and $\mathcal{L}_{spread}$ of visual representation domain [29, 30] in the Appendix. Our $\mathcal{L}_{spread}$ is formulated as: $\mathcal{L}_{spread}=\frac{1}{2}\left[\mathcal{L}_{spread-\text{VL}}+\mathcal{L}_{\text{LV}}\right]$ where $\mathcal{L}_{spread-\text{VL}}=(1-\alpha)\mathcal{L}_{ConCon}+\alpha\mathcal{L}_{contextNCE}$ and $\mathcal{L}_{\text{LV}}$ is from (1).
Spread Loss – Vision-to-Language (VL): $\mathcal{L}_{spread-\text{VL}}$
@@
The final formulation for VL direction is a weighted sum: $\mathcal{L}_{spread-\text{VL}}=(1-\alpha)\mathcal{L}_{ConCon}+\alpha\mathcal{L}_{contextNCE}$, where $\alpha$ controls the trade-off between context-level alignment and intra-context contrast. While $\mathcal{L}_{ConCon}$ allows similar context to be close, $\mathcal{L}_{contextNCE}$ prevents representation collapse in similar context embeddings. Thus, $\mathcal{L}_{spread}$ enables learning representations that are not only globally aligned across modalities but also sensitive to subtle mismatches in local content. In our experiments, we use $\alpha=0.5$ to balance the two VL components.
Spread Loss – Language-to-Vision (LV): $\mathcal{L}_{\text{LV}}$
Since we aim to preserve bidirectional learning throughout the joint-AE training, for the LV direction, we align positive text to its corresponding image and penalize non-matching texts (excluding hard-negative texts since there are no matching images) using a standard contrastive approach. Hence, it stays the same as $\mathcal{L}_{\text{LV}}$ used in $\mathcal{L}_{Con}$ and $\mathcal{L}_{NegCon}$.
Language Backbone Vision Backbone Replace Task Add Task Swap Task
(Model Size) (Model Size) Method Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way)
Gemma2 (2B) DINOv2 (ViT-B; 86M) Con 65.14 59.97 58.34 54.20 57.36 51.005
NegCon 86.32 81.05 96.85 95.28 74.5 66.30
Spread 88.01 82.32 95.74 93.14 80.2 70
Llama3.2 (1B) DINOv2 (ViT-B) Con 66.97 60.93 60.99 59.42 66.59 58.36
NegCon 83.43 76.73 94.25 92.17 68.26 64.29
Spread 87.21 83.61 97.83 91.57 79.05 66.57
OLMo2 (7B) DINOv2 (ViT-B) Con 68.48 65.73 65.17 62.22 66.88 62.91
NegCon 85.30 78.13 96.67 97.96 71.69 66.88
Spread 88.32 80.10 97.41 92.16 77.62 70.27
Pretrained CLIP (ViT-B-32/OpenAI) [32] 81.05 80.40 77.58 76.95 64.69 64.41
Pretrained CLIP (ViT-B-32/LAION-400m) [32, 33] 80.90 80.28 79.64 79.25 67.30 67.22
Pretrained CLIP (ViT-B-32/SigLIP) [34] 85.01 83.80 86.56 85.97 70.97 69.91
Finetuned CLIP 74.62 72.38 92.69 87.4 67.21 65.4
Table 2: Image-to-Text Retrieval Results of Joint Autoencoder Modulator (JAM) for Vision-Language Compositionality. The colored values match the backbones analyzed in Figure 1. Replace, Add and Swap tasks are described in Section 2.1. We report Recall@1 scores (binary and 5-way) across three compositional tasks: Replace, Add, and Swap, using models trained with different alignment methods (Con, NegCon, Spread). JAM with Spread loss consistently outperforms contrastive baselines across all tasks and backbones. Morever, it matches or surpasses several strong pretrained and finetuned CLIP variants, highlighting the effectiveness of structured alignment over independently pretrained representations for fine-grained vision–language reasoning.
Language Backbone Vision Backbone Replace Task Add Task Swap Task
(Model Size) (Model Size) Method Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way) Recall@1 (binary) Recall@1 (5-way)
NegCon 87.52 84.33 98.89 93.14 84.44 69.69
Spread 89.66 83.34 98.89 96.30 84.44 73.38
Llama3.2 (3B) DINOv2 (ViT-L) Con 65.24 59.14 62.38 57.9 73.12 65.18
NegCon 84.54 82.66 92.61 88.16 75.93 70.27
Spread 89.43 82.48 95.74 91.25 82.44 76.77
OLMo2 (13B) DINOv2 (ViT-L) Con 65.86 68.68 59.21 62.52 62.28 52.12
NegCon 89.36 82.68 93.7 91.66 78.89 67.57
Spread 90.53 83.18 97.96 92.42 84.71 69.69
Table 3: Model Scaling Experiment We report image-to-text Recall@1 scores across three spatial reasoning tasks (Replace, Add, Swap) using language backbones of increasing scale—Gemma2 (9B), LLaMA 3.2 (3B), and OLMo2 (13B)—within our JAM framework. Despite varying parameter counts, we observe no consistent trend of performance improvement with model size. Spread loss performs robustly across all scales, suggesting that for fine-grained, low-data tasks, representational alignment depends more on objective design than on model scale.

### v5

@@ Abstract
Independently trained vision and language models inhabit disjoint representational spaces, shaped by their respective modalities, objectives, and architectures. The Platonic Representation Hypothesis (PRH) suggests these models may nonetheless converge toward a shared statistical model of reality. This raises a fundamental question: can we move beyond post-hoc detection of such alignment and explicitly optimize for it? We argue this challenge is most critical in fine-grained contextual distinctions—where multiple descriptions share global semantics but differ in subtle compositional details. We address this with the Joint Autoencoder Modulator (JAM), which aligns frozen unimodal models by jointly training modality-specific autoencoders with coordinated reconstruction and cross-modal alignment objectives. We systematically evaluate JAM across three design axes: (i) alignment objectives, introducing our multimodal Spread Loss that outperforms classic contrastive methods; (ii) the layer depth at which alignment is most effective; and (iii) the role of foundation model scale in representational convergence. Our findings show that JAM reliably induces alignment even across independently trained representations, offering both theoretical insight into the structure of shared semantics and practical guidance for transforming generalist unimodal foundations into specialist multimodal models.
@@ 1 Introduction
Neural networks trained on different modalities, datasets, and objectives typically inhabit disjoint representational spaces. Yet the Platonic Representation Hypothesis (PRH) [1] suggests that these models—despite having no shared supervision, architecture, or training regime—may nonetheless converge toward a common statistical model of reality. The naming of PRH is inspired by Plato’s Allegory of the Cave [2], which describes individuals who observe only shadows of objects cast on the wall of a cave and mistake these projections for the entirety of reality. This hypothesis has been discussed under several philosophical and empirical lenses, including convergent realism in the philosophy of science and the Anna Karenina scenario [3] in representation learning, which suggests that all well-performing models may ultimately resemble each other.
However, most evidence remain observational, relying on statistical metrics that infer compatibility at a coarse-grained level. These methods quantify global correlations across feature spaces, but do not provide practical mechanisms for constructing multimodal systems from unimodal ones, nor do they illuminate where alignment fails. Crucially, the limitations emerge in fine-grained contextual settings, where multiple candidate descriptions may share overall semantics yet differ in specific details. For example, distinguishing whether an image contains a dog is a coarse judgment, but deciding between “a brown dog chasing a red ball” and “a brown dog chasing a blue ball” requires contextual sensitivity to subtle compositional cues. It is precisely in these cases (i.e., where local semantic differences matter) that the true challenge of alignment lies.
1.1 Formulation of Plato’s Cave
We cast Plato’s Cave in two regards:
  • Unimodal representations: Borrowing from the Plato’s Cave allegory, each unimodal data and training regime yields only a “shadowed projection” of reality. Our goal is to uncover the latent, Platonic representation: a shared semantic structure that exists despite the models being trained in isolation.
  • Fine-grained context: While unimodal embeddings capture coarse contextual structure (e.g., broad semantic overlap), they often fail to encode discriminative fine-grained details. Here, “context” refers to the high-level semantics shared across modalities, while “fine-grained” refers to resolving distinctions within that shared context (e.g., attributes, relations, or localized compositional shifts).
@@ 1.2 Escaping Plato’s Cave (Platonic Alignment)
Motivated by these insights, we propose Platonic Alignment, a framework to explicitly align unimodal models trained independently on distinct modalities. Our Joint Autoencoder Modulator (JAM, Fig. 3) aligns frozen language and vision representations using coordinated reconstruction and alignment objectives. Reconstruction preserves modality-specific information, while a shared bottleneck enforces a coherent conceptual space capable of resolving both coarse and fine-grained semantics.
Our contributions are as follows:
  • We introduce the multimodal Spread Loss, which leverages contextual structure to outperform classic contrastive objectives for fine-grained alignment.
  • We demonstrate the versatility of our JAM framework across a wide range of pretrained backbones, and validate its components through ablations and comparisons to projection-based baselines.
  • We analyze the impact of pretrained backbone/model’s scale and layer depth on alignment performance, providing insights into how these factors interact with alignment supervision.
@@ 1.3 Related Work
Testing for Alignment. Prior work in this direction has been largely diagnostic focusing on evaluating alignment between frozen features with broad, context-agnostic datasets (e.g., Wikipedia caption dataset (WIT) [4]): measures such as centered kernel alignment (CKA) [5], variants of CCA (SVCCA [6], projection-weighted CCA [7]), and nearest-neighbor metrics [8, 1]. Other approaches explore probing tasks or zero-shot transfer to assess latent compatibility across modalities [9, 10]. More recently, model-stitching frameworks have examined whether independently trained sub-networks can be functionally composed to perform new tasks [3], hinting at deeper interoperability between pretrained systems but focusing on vision domain [3, 11].
@@ 2 Statistical Tests for Representation Alignment
To probe the potential for aligning unimodal models, we applied four alignment metrics—CCA, CKA, SVCCA, and CKNNA—to three types of image–text pairs, illustrated in Fig. 1: match (true positive, image–caption pairs), easy non-match (unrelated captions), and hard non-match (semantically similar captions differing in fine-grained details). The hard negatives are particularly vital in our setting: While they share global semantics with the true captions, they diverge in localized compositional attributes, testing whether alignment captures fine-grained context rather than coarse similarity.
Figure 1: Illustration of fine-grained contextual understanding from the SugarCrepe dataset [20]. Each image is paired with three types of captions: (i) Match (true positive) captions that correctly describe the image, (ii) Easy non-match captions that are entirely unrelated, and (iii) Hard non-match (hard negative) captions that share global semantics with the true caption but diverge in subtle, fine-grained details (e.g., swapping relations, replacing objects, or adding attributes). These controlled perturbations—Replace (R), Swap (S), and Add (A) from the SugarCrepe dataset [20] provide a principled test for whether alignment methods capture contextual fine-grained distinctions rather than just coarse semantic similarity.
We evaluate embeddings from unimodally pretrained vision and language models (Gemma2 (2B) [21], Llama3.2 (1B) [14], OLMo2 (7B) [22] for language; DINOv2 (ViT-B) [23], ResNet50 [24] for vision), using the SugarCrepe dataset [20], which is explicitly designed for fine-grained vision–language compositionality. SugarCrepe provides minimal caption perturbations across three transformation families—Replace, Swap, and Add—yielding controlled hard negatives (see Fig. 1). For experiments, we extract CLS-token image embeddings and penultimate-layer text embeddings, then compute alignment scores using the four metrics (See Appendix for metrics formulations).
@@
Figure 2: Statistical Metrics for Representation Alignment: Across all metrics and model cases, match pairs consistently show higher alignment scores than easy non-match pairs, supporting the hypothesis that unimodal models encode shared global structure. However, hard non-match pairs exhibit similarly high scores. This indicates that while statistical metrics for representations reveal coarse representational compatibility – which aligns with PRH [1] – but are insufficient for diagnosing semantic alignment at a more granular level. (a) Linear and Kernel (Gaussian (i.e., RBF) kernel) PCA to reduce the embeddings dimension to 50; Set CCA dimension as 50. Variation of score values are based on the utilized kernel. (b) Variation of score values are based on the utilized kernel (Linear & RBF). (c) SVD to reduce embeddings dimension to 10; Set CCA dimension as 10. (d) Set top k nearest neighbors = 10 (following the default setting of [1]).
@@
With the extracted feature representations  [^1] All experiments were conducted with non-GDM (Google DeepMind) compute resources., let $D$ be the whole data, and $B$ be the batch. We construct $D$ in the nested pair format: $D=\{(v_{i},(l_{P_{i}},l_{N_{i}})\}_{i=1}^{n}$ (Refer to Table 6 for variable descriptions). $C(i)$, the set of similar context text embedding for anchor $i$, is defined as $l_{P_{i}}$ and its hard negative $l_{N_{i}}$ (i.e., $C(i)=\{l_{P_{i}},l_{N_{i}}\}$). Accordingly, $\widetilde{C}(i)=L\setminus C(i)$. (In Table 2, we show results for DINOv2 as vision backbone. For full results on refer to Appendix Table 7,8,9.)
2.1 Statistical tests detect alignment, but not fine-grained context.
@@
The results (Fig.2) show that all models reliably distinguished matching from easy non-matching pairs, confirming that independently trained vision and language models encode a shared high-level structure consistent with the Platonic Representation Hypothesis [1]. However, in the hard non-match condition, statistical metrics produced scores nearly indistinguishable from true matches. This is not a failure of alignment but a reflection of the metrics themselves: hard negatives share global semantics with positives, differing only in fine-grained attributes. We conjecture that representational similarity is real at the coarse level but insufficiently sensitive to discriminative detail. Thus, while statistical tests reveal broad compatibility, they obscure nuanced distinctions critical for fine-grained alignment. To move beyond this limitation, we turn to a lightweight post-hoc training approach that explicitly optimizes for semantic coherence, both globally and locally. This insight directly led to our alignment training scheme in the next section.
@@ 3 Our method: Joint Autoencoder Modulator (JAM)
3.1 Loss functions
For the objective design (i.e., loss functions) to jointly train autoencoders for alignment (along with default reconstruction objective), we adopt Contrastive and Negative Contrastive loss, which are utilized as baseline methods, and further introduce our new loss scheme: Spread Loss functions for vision-to-language and language-to-vision.
Having a set of losses designed to optimize different aspects of the alignment had a side benefit of helping us to diagnose and analyze experiments (e.g., which layer to establish alignment), as will be shown in later sections.
Contrastive Loss (Con)
@@ 3.1 Loss functions
To address this, we introduce $\mathcal{L}_{spread}$, a contrastive objective that incorporates a notion of context similarity and fine-grained differentiation. As shown in Figure 4, the core idea is to pull together all semantically similar captions, including the ground-truth caption and selected hard negatives, and then selectively push apart fine-grained distinctions within that similar group. This two-stage formulation allows us to preserve shared contextual alignment while enforcing discriminative power at the instance level. The idea of spread loss has been proposed in the domain of visual representation learning [31, 32], which is a variant of Supervised Contrastive (SupCon) loss [33], but tailored to tackling class collapse. We take inspiration from that work and apply this idea to the multi-modal alignment problem. [^2] We provide a parallel comparison between $\mathcal{L}_{spread}$ of multimodal representation space (i.e., ours) and $\mathcal{L}_{spread}$ of visual representation domain [31, 32] in the Appendix. Our $\mathcal{L}_{spread}$ is formulated as:
where $\mathcal{L}_{spread-\text{VL}}=(1-\alpha)\mathcal{L}_{ConCon}+\alpha\mathcal{L}_{contextNCE}$ and $\mathcal{L}_{\text{LV}}$ is from (1), preserving bidirectional learning throughout joint-AE training.
@@
The final formulation for VL direction is a weighted sum: $\mathcal{L}_{spread-\text{VL}}=(1-\alpha)\mathcal{L}_{ConCon}+\alpha\mathcal{L}_{contextNCE}$, where $\alpha$ controls the trade-off between context-level alignment and intra-context contrast. While $\mathcal{L}_{ConCon}$ allows similar context to be close, $\mathcal{L}_{contextNCE}$ prevents representation collapse in similar context embeddings. Thus, $\mathcal{L}_{spread}$ enables learning representations that are not only globally aligned across modalities but also sensitive to subtle mismatches in local content. In our experiments, we use $\alpha=0.5$ to balance the two components in $\mathcal{L}_{spread-\text{VL}}$.